Circular No. 7208

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1999 L5 (SOHO) AND C/1999 M1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports his detection and measurements (reduced by the undersigned) of two more apparent Kreutz sungrazing comets in images obtained with both the C2 and C3 coronagraphs aboard SOHO. Neither comet showed evidence of a tail. Further details are given on MPEC 1999-M23 and 1999-M24.

1999 UT  α_{2000}  δ_{2000}  Comet
June 10.613  5^{h}10.^{m}1  +20^{o}13′  C/1999 L5
16.138  5 32.8  +20 49  C/1999 M1
1999 LD_{31} AND 1999 LE_{31}

MPEC 1999-M28 and 1999-M29 provide detailed information about two apparently asteroidal objects, 1999 LD_{31} and LE_{31}, discovered by LIN-EAR on June 8 and 12, respectively (with prediscovery observations of the latter on May 17), and followed extensively by observers using The NEO Confirmation Page. In each case the orbit has been found to be retrograde: 1999 LD_{31} has a = 21.9 AU, e = 0.89, i = 160^{o}, H = 13.9; 1999 LE_{31} has a = 8.0 AU, e = 0.46, i = 152^{o}, H = 12.3. All observers have consistently reported 1999 LD_{31} to be asteroidal, and only one observer has suggested that 1999 LE_{31} may have cometary appearance (although this is unconfirmed). In particular, A. Fitzsimmons, Queen’s University, Belfast, reports that 250-s exposures in 1′′ seeing by S. Collander-Brown and S. Lowry with the 1-m Kapteyn telescope at La Palma on June 15 show both objects clearly to be point sources.

NOVA IN NGC 6822 J. Y. King and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), report the discovery of an apparent nova in NGC 6822, a dwarf irregular galaxy that is probably in the Local Group. The nova was discovered on unfiltered images taken on June 23.40 and 23.44 UT (mag ∼ 17.3) and confirmed by an observation made on June 24.38 (mag ∼ 17.0). The new object is located at α = 19^{h}45^{m}00^{s}.31, δ = −14^{o}50′10′′.3 (equinox 2000.0). A KAIT image of the same field on June 19.4 UT (limiting magnitude ∼ 18.5) showed nothing at the position of the nova. A finding chart for the nova can be

found at the URL http://astron.berkeley.edu/~bait/n6822.gif.”

1999 June 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden